# Online Price Competition under Generalized Linear Demands

### v3

@@ 2.1 Regret and Nash equilibrium
$\Gamma_{i}(\mathbf{p}_{-i})\triangleq\underset{p_{i}\in\mathcal{P}_{i}}{\arg\max}\operatorname{rev}_{i}(p_{i}|\mathbf{p}_{-i}),$ (4)
@@
2.8 is common in the literature (bracale2025revenue; scutari2014real; li2024lego; tsekrekos2024variational) and includes linear demand models (li2024lego), concave demand specifications, and $s_{i}$-concave demand functions with $s_{i}\geq 1$ (see bracale2025revenue, Appendix I).
2.8 guarantees that the best-response map $\Gamma_{i}(\cdot)$ in (4) is well defined, i.e., admits a unique value for every $\mathbf{p}_{-i}$, which can be found by solving the first-order condition $\partial_{p_{i}}\operatorname{rev}_{i}(p_{i}|\mathbf{p}_{-i})=0$, w.r.t. $p_{i}$ and projecting the solution on $\mathcal{P}_{i}$. Additionally, 2.8 guarantees that the learning dynamics converge with sublinear sellers’ regrets. Indeed, as in li2024lego; bracale2025revenue, we control the individual regret upper bound by controlling the $L^{2}$ difference of the joint policies with respect to the NE, i.e. $\operatorname{Reg}_{i}(T)\lesssim\sum_{t=1}^{T}\|\mathbf{p}^{(t)}-\mathbf{p}^{*}\|_{2}^{2}$ where $\mathbf{p}^{*}$ is a NE (see Equation 15 for a detailed bound). A Nash equilibrium $\mathbf{p}^{*}=\left(p_{i}^{*}\right)_{i\in[N]}\in\mathcal{P}$ is defined as a price vector under which unilateral deviation is not profitable for any seller. Specifically,
@@
where $\boldsymbol{\Gamma}:\mathcal{P}\rightarrow\mathcal{P}$ is called best-response operator. The next result is an immediate consequence of Theorem 3 in scutari2014real.
@@
The log-concavity assumption covers the linear demand model by li2024lego for which $\mu_{i}(u)=\alpha_{i,0}+u$ is log-concave. This assumption can be found in several works with $N=1$ (monopolistic setting); for example, Assumption 2.1 in fan2024policy, Assumption 1 in chen2018robust, Equation 1 in cole2014sample and golrezaei2019dynamic; javanmard2017perishability; javanmard2019dynamic which assume $\mu_{i}$ and $1-\mu_{i}$ to be log-concave, where $\mu_{i}$ is a survival function. The assumption in the model parameter constraint coincides with Assumption 2 in li2024lego, and similar assumptions can be found in bracale2025revenue; kachani2007modeling.
@@
Lemma 2.11 was proved by bracale2025revenue in the more general case of $s_{i}$-concave demand model for $s_{i}>-1$ (log-concavity is covered for $s_{i}=0$) and implies that the NE is unique by the Banach Contraction theorem. As in li2024lego; bracale2025dynamic the uniqueness of the NE entails sublinear individual regret for all sellers, as we will see in the proof of our main Theorem 4.1.
@@ 3 Algorithm
We adopt a similar algorithm to russac2020algorithms based on UCB of the revenue functions $\operatorname{rev}_{i}(\cdot|\mathbf{p}_{-i})$ for every $\mathbf{p}_{-i}\in\mathcal{P}_{-i}$. This method relies on a penalized log-likelihood approach that has a regularizing effect and avoids the need for common initial exploration across the sellers.
@@ 3.1 Penalized MLE
where the LHS is the derivative of the penalized log-likelihood in (6). Following the notation in russac2020algorithms, we introduce
@@ 3.2 Pricing Policy via UCB
Following standard UCB-based approaches (russac2020algorithms; russac2019weighted), we define the upper confidence bound (UCB) of the estimated revenue as
@@
[…]
  if $\lVert\widehat{\boldsymbol{\theta}}_{i}^{(t)\tt PML}\rVert_{2}\leq B_{\theta_{i}}$ let $\widetilde{\boldsymbol{\theta}}_{i}^{(t)\tt PML}=\widehat{\boldsymbol{\theta}}_{i}^{(t)\tt PML}$ else compute $\widetilde{\boldsymbol{\theta}}_{i}^{(t)\tt PML}$ defined in (9).
[…]
@@ 4 Regret Analysis
At round $t$, seller $i$ updates the matrix $V_{i}^{(t)}=V_{i}^{(t-1)}+\mathbf{p}^{(t)}\mathbf{p}^{(t)\top}$ and its inverse. Using the Sherman–Morrison formula to update $V_{i}^{(t),-1}$ costs $O(N^{2})$ time (and forming $V_{i}^{(t)}$ itself is also $O(N^{2})$). To evaluate the UCB bonus for a candidate price $p_{i}$, one needs $\|(p_{i},\mathbf{p}_{-i})\|_{V_{i}^{(t),-1}}$, which requires a matrix–vector product $V_{i}^{(t),-1}\mathbf{p}$ in $O(N^{2})$ time and then a dot product in $O(N)$. If the price is chosen by evaluating a grid of $G$ candidate prices (or an equivalent one-dimensional search that uses $G$ evaluations), the per-round cost per seller is $O(GN^{2})$. Over $T$ rounds, the time is $O(TGN^{2})$ per seller, and $O(TGN^{3})$ if all $N$ sellers are run. Space usage per seller is $O(N^{2})$ to store $V_{i}^{(t)}$ and (optionally) $V_{i}^{(t),-1}$ plus $O(N)$ for current vectors; if past histories are not stored, $O(N^{2})$ dominates.
@@ 4.1 Convergence to Nash Equilibrium
The bound in (13) alone does not imply convergence of $\mathbb{E}[\|\mathbf{p}^{(T)}-\mathbf{p}^{*}\|_{2}^{2}]$. This happens of course if $\mathbb{E}[\|\mathbf{p}^{(t)}-\mathbf{p}^{*}\|_{2}^{2}]$ is monotonically decreasing in $t$, as in li2024lego and bracale2025revenue due to the initial i.i.d. pricing exploration in their algorithm. Under such an i.i.d. exploration phase, those works establish that $\|\widehat{\boldsymbol{\theta}}_{i}^{(t)}-\boldsymbol{\theta}_{i,0}\|_{2}=O_{P}(t^{-1/2})$ under standard regularity conditions, which in turn implies $\mathbb{E}[\|\mathbf{p}^{(T)}-\mathbf{p}^{*}\|_{2}^{2}]=O(T^{-1/2})$. Unfortunately, this is not the case in our setting due to the absence of i.i.d. pricing estimation, since the prices $\{\mathbf{p}^{(t)}\}_{t\in[T]}$ are inherently adaptive rather than i.i.d., as their selection depends on past noise realizations. In the absence of i.i.d. pricing excitation, estimation error can only be controlled along the directions that are sufficiently explored. As a result, $\|\widehat{\boldsymbol{\theta}}_{i}^{(t)}-\boldsymbol{\theta}_{i,0}\|_{2}$ does not generally decay at the parametric rate $O_{P}(t^{-1/2})$. Nevertheless, as formally established by our Theorem 4.1, achieving optimal (or near-optimal) regret remains possible, which is consistent with the general bandit literature, where the same phenomenon is observed – for example, in OFU and UCB-type algorithms for linear and generalized linear bandits (dani2008stochastic; abbasi2011improved; filippi2010parametric; russac2019weighted).
5 Conclusions
Limitations.
@@ 5 Conclusions
We assume a known increasing link $\mu_{i}$ with mild smoothness/curvature, bounded prices and compact parameter sets, and strong concavity of per-round revenues together with a contraction-type condition that guarantees a unique NE. None of these is new: they are standard in dynamic pricing and learning-in-games and strictly include the linear-demand setting of li2024lego as a special case. These conditions ensure well-posed best responses and well-conditioned information for concentration, but they also circumscribe our scope to stationary parameters and synchronous price moves under a unique-equilibrium regime.
Future directions.
@@
A natural avenue is to relax link knowledge by replacing it with shape constraints (e.g., monotonicity or $s$-concavity) and to derive calibration-free confidence bonuses that adapt to local curvature while maintaining a non-cooperative initial exploration phase. On the game-theoretic side, moving beyond global contractions – so as to allow multiple or set-valued equilibria – suggests a variational-inequality/monotone-operator treatment to obtain no-regret convergence. Algorithmically, Bayesian or Thompson sampling approaches to GLMs under our information structure (where rivals’ revenues remain unobserved) are particularly appealing, as they tend to yield less conservative confidence bounds in practice compared to UCB-based methods. Finally, incorporating nonstationarity in $\boldsymbol{\theta}_{i,0}$, inventory constraints, and asynchronous observability of rivals’ prices would bring the model even closer to more real market scenarios.
[…]
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes]
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes: Remark 4.4]
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Not Applicable]
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes]
    2. (b)
      Complete proofs of all theoretical results. [Yes]
    3. (c)
      Clear explanations of any assumptions. [Yes]
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Not Applicable]
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Not Applicable]
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Not Applicable]
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Not Applicable]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Not Applicable]
    2. (b)
      The license information of the assets, if applicable. [/Not Applicable]
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Not Applicable]
    4. (d)
      Information about consent from data providers/curators. [Not Applicable]
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Not Applicable]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]
Supplementary for
Online Price Competition under Generalized Linear Demands
@@
For the first part of the proof, we follow similar steps to russac2020algorithms, Corollary 1. We start by upper-bounding the following difference:
@@
For simplicity of notation, we write $\widehat{\boldsymbol{\theta}}_{i}^{(t)}$ for $\widehat{\boldsymbol{\theta}}_{i}^{(t)\tt PML}$ and $\widetilde{\boldsymbol{\theta}}_{i}^{(t)}$ for $\widetilde{\boldsymbol{\theta}}_{i}^{(t)\tt PML}$. This proof follows a similar structure to the proof of Proposition 1 in russac2020algorithms. Recall the definitions
@@
In the last inequality, we have used the assumption that $\lVert\boldsymbol{\theta}_{i,0}\rVert_{2}\leq B_{\theta_{i}}$ and the concentration result established in Proposition 7 of russac2019weighted for the self-normalized quantity $\|S_{i}^{(t-1)}\|_{V_{i}^{(t-1),-1}}$ (which is applicable thanks to the conditional $L_{\mu_{i}}$-subgaussianity of $\eta_{i}^{(t)}$ established in Lemma 2.6), which establishes that
@@
We use the Elliptical Potential Lemma in carpentier2020elliptical Proposition 1 (by replacing $p\leftarrow 1$ and $\lambda\leftarrow\frac{\lambda}{c_{\mu_{j}}}$):

### v4

@@ 2.1 Regret and Nash equilibrium
$\Gamma_{i}(\mathbf{p}_{-i})\triangleq\arg\max_{p_{i}\in\mathcal{P}_{i}}\,\operatorname{rev}_{i}(p_{i}|\mathbf{p}_{-i}),$ (4)
@@
2.8 is common in the literature (Bracale et al., 2025a; Scutari et al., 2014; Li et al., 2024; Tsekrekos and Yannacopoulos, 2024) and includes linear demand models (Guda and Subramanian, 2026; Li et al., 2024), concave demand specifications, and $s_{i}$-concave demand functions with $s_{i}\geq 1$ (see Bracale et al. (2025a), Appendix I).
2.8 guarantees that the best-response map $\Gamma_{i}(\cdot)$ in (4) is well defined, i.e., admits a unique value for every $\mathbf{p}_{-i}$, which can be found by solving the first-order condition $\partial_{p_{i}}\operatorname{rev}_{i}(p_{i}|\mathbf{p}_{-i})=0$, w.r.t. $p_{i}$ and projecting the solution on $\mathcal{P}_{i}$. Additionally, 2.8 guarantees that the learning dynamics converge with sublinear sellers’ regrets. Indeed, as in Li et al. (2024); Bracale et al. (2025a), we control the individual regret upper bound by controlling the $L^{2}$ difference of the joint policies with respect to the NE, i.e. $\operatorname{Reg}_{i}(T)\lesssim\sum_{t=1}^{T}\|\mathbf{p}^{(t)}-\mathbf{p}^{*}\|_{2}^{2}$ where $\mathbf{p}^{*}$ is a NE (see Equation 15 for a detailed bound). A Nash equilibrium $\mathbf{p}^{*}=\left(p_{i}^{*}\right)_{i\in[N]}\in\mathcal{P}$ is defined as a price vector under which unilateral deviation is not profitable for any seller. Specifically,
@@
where $\boldsymbol{\Gamma}:\mathcal{P}\rightarrow\mathcal{P}$ is called best-response operator. The next result is an immediate consequence of Theorem 3 in Scutari et al. (2014).
@@
The log-concavity assumption covers the linear demand model by Li et al. (2024) for which $\mu_{i}(u)=\alpha_{i,0}+u$ is log-concave. This assumption can be found in several works with $N=1$ (monopolistic setting); for example, Assumption 2.1 in Fan et al. (2024), Assumption 1 in Chen and Farias (2018), Equation 1 in Cole and Roughgarden (2014) and Golrezaei et al. (2019); Javanmard (2017); Javanmard and Nazerzadeh (2019) which assume $\mu_{i}$ and $1-\mu_{i}$ to be log-concave, where $\mu_{i}$ is a survival function. The assumption in the model parameter constraint coincides with Assumption 2 in Li et al. (2024), and similar assumptions can be found in Bracale et al. (2025a); Kachani et al. (2007).
@@
Lemma 2.11 was proved by Bracale et al. (2025a) in the more general case of $s_{i}$-concave demand model for $s_{i}>-1$ (log-concavity is covered for $s_{i}=0$) and implies that the NE is unique by the Banach Contraction theorem. As in Li et al. (2024); Bracale et al. (2025b) the uniqueness of the NE entails sublinear individual regret for all sellers, as we will see in the proof of our main Theorem 4.1.
@@ 3 Algorithm
We adopt a similar algorithm to Russac et al. (2020) based on UCB of the revenue functions $\operatorname{rev}_{i}(\cdot|\mathbf{p}_{-i})$ for every $\mathbf{p}_{-i}\in\mathcal{P}_{-i}$. This method relies on a penalized log-likelihood approach that has a regularizing effect and avoids the need for common initial exploration across the sellers.
@@ 3.1 Penalized MLE
where the LHS is the derivative of the penalized log-likelihood in (6). Following the notation in Russac et al. (2020), we introduce
@@ 3.2 Pricing Policy via UCB
Following standard UCB-based approaches (Russac et al., 2020, 2019), we define the upper confidence bound (UCB) of the estimated revenue as
@@
  $\bullet$ Each seller $i$ computes $\widehat{\boldsymbol{\theta}}_{i}^{(t)\tt PML}$ according to (7). If $\lVert\widehat{\boldsymbol{\theta}}_{i}^{(t)\tt PML}\rVert_{2}\leq B_{\theta_{i}}$ let $\widetilde{\boldsymbol{\theta}}_{i}^{(t)\tt PML}=\widehat{\boldsymbol{\theta}}_{i}^{(t)\tt PML}$. Else compute $\widetilde{\boldsymbol{\theta}}_{i}^{(t)\tt PML}$ defined in (9).
  $\bullet$ Each seller $i$ set $p_{i}^{(t)}\leftarrow\widehat{\Gamma}_{i}^{(t)}(\mathbf{p}^{(t-1)}_{-i})$, defined in (12).
  $\bullet$ Each seller $i$ observes $y_{i}^{(t)}$.
  $\bullet$ Each seller $i$ observes $\mathbf{p}^{(t)}$.
  $\bullet$ Each seller $i$ updates $V_{i}^{(t)}\leftarrow V_{i}^{(t-1)}+\mathbf{p}^{(t)}\mathbf{p}^{(t)\top}$.
@@ 4 Regret Analysis
At round $t$, seller $i$ updates the matrix $V_{i}^{(t)}=V_{i}^{(t-1)}+\mathbf{p}^{(t)}\mathbf{p}^{(t)\top}$ and its inverse. Using the Sherman–Morrison formula to update $V_{i}^{(t),-1}$ costs $O(N^{2})$ time (and forming $V_{i}^{(t)}$ itself is also $O(N^{2})$). To evaluate the UCB bonus for a candidate price $p_{i}$, one needs $\|(p_{i},\mathbf{p}_{-i})\|_{V_{i}^{(t),-1}}$, which requires a matrix–vector product $V_{i}^{(t),-1}\mathbf{p}$ in $O(N^{2})$ time and then a dot product in $O(N)$. Both the true best response and the optimistic best response solve a one-dimensional concave maximization problem. In practice, this can be implemented via a 1-D convex optimization (e.g., bisection or Newton), which is costless, i.e., $O(1)$ (this implies that the UCB optimization is tractable and does not suffer from the NP-hardness of high-dimensional problems). Over $T$ rounds, the time is $O(TN^{2})$ per seller, and $O(TN^{3})$ if all $N$ sellers are run. Space usage per seller is $O(N^{2})$ to store $V_{i}^{(t)}$ and (optionally) $V_{i}^{(t),-1}$ plus $O(N)$ for current vectors; if past histories are not stored, $O(N^{2})$ dominates.
@@ 4.1 Convergence to Nash Equilibrium
The bound in (13) alone does not imply a convergence rate for $\mathbb{E}\left[\|\mathbf{p}^{(T)}-\mathbf{p}^{*}\|_{2}^{2}\right]$ as $T\to\infty$. Indeed, a bound on the partial sums only yields pointwise convergence if the sequence is sufficiently regular (e.g., eventually monotone), or under an initial i.i.d. pricing exploration phase, like in Li et al. (2024) and Bracale et al. (2025a), which ensures that the estimation error satisfies $\|\widehat{\boldsymbol{\theta}}_{i}^{(T)}-\boldsymbol{\theta}_{i,0}\|_{2}=O_{P}(T^{-1/2})$ under standard regularity conditions, which in turn yields the parametric-rate convergence $\mathbb{E}\left[\|\mathbf{p}^{(T)}-\mathbf{p}^{*}\|_{2}^{2}\right]=O(T^{-1/2})$.
In our setting, however, prices $\{\mathbf{p}^{(t)}\}_{t\leq T}$ are generated adaptively (their selection depends on past noise realizations), and we do not impose an i.i.d. excitation phase. In the absence of this phase, the estimation error can typically be controlled only along directions that are sufficiently excited by the played actions. As a result, $\|\widehat{\boldsymbol{\theta}}_{i}^{(T)}-\boldsymbol{\theta}_{i,0}\|_{2}$ does not generally decay at the parametric rate $O_{P}(T^{-1/2})$, and one should not expect a pointwise convergence rate for $\mathbb{E}\left[\|\mathbf{p}^{(T)}-\mathbf{p}^{*}\|_{2}^{2}\right]$ based solely on (13).
Nevertheless, as we show in Theorem 4.1, our primary objective is regret optimality, which can be achieved even without strong pointwise convergence of the price sequence to a Nash equilibrium. This is consistent with the broader bandit literature, where regret-optimal algorithms often rely on adaptive (non-i.i.d.) action selection, as in OFU/UCB-style methods for linear and generalized linear bandits (Dani et al., 2008; Abbasi-Yadkori et al., 2011; Filippi et al., 2010; Russac et al., 2019).
5 Limitations and future directions.
@@
We assume a known increasing link $\mu_{i}$, strong concavity of per-round revenues, and a contraction-type condition that guarantees a unique NE. None of these is new: they are standard in dynamic pricing and learning-in-games and strictly include the linear-demand setting of Li et al. (2024) as a special case. These conditions ensure well-posed best responses and well-conditioned information for concentration, but they also circumscribe our scope to stationary parameters and synchronous price moves under a unique-equilibrium regime.
@@
A natural avenue is to relax link knowledge by replacing it with shape constraints (e.g., monotonicity or $s$-concavity) and to derive calibration-free confidence bonuses that adapt to local curvature while remaining free of any initial exploration phase. On the game-theoretic side, moving beyond global contractions – to allow multiple or set-valued equilibria – suggests a variational-inequality/monotone-operator treatment to obtain no-regret convergence. Algorithmically, Bayesian or Thompson sampling approaches to GLMs under our information structure (where rivals’ revenues remain unobserved) are particularly appealing, as they tend to yield less conservative confidence bounds in practice compared to UCB-based methods. Finally, incorporating nonstationarity in $\boldsymbol{\theta}_{i,0}$, inventory constraints, and asynchronous observability of rivals’ prices would bring the model even closer to more real market scenarios.
@@
For the first part of the proof, we follow similar steps to Russac et al. (2020), Corollary 1. We start by upper-bounding the following difference:
@@
For simplicity of notation, we write $\widehat{\boldsymbol{\theta}}_{i}^{(t)}$ for $\widehat{\boldsymbol{\theta}}_{i}^{(t)\tt PML}$ and $\widetilde{\boldsymbol{\theta}}_{i}^{(t)}$ for $\widetilde{\boldsymbol{\theta}}_{i}^{(t)\tt PML}$. This proof follows a similar structure to the proof of Proposition 1 in Russac et al. (2020). Recall the definitions
@@
In the last inequality, we have used the assumption that $\lVert\boldsymbol{\theta}_{i,0}\rVert_{2}\leq B_{\theta_{i}}$ and the concentration result established in Proposition 7 of Russac et al. (2019) for the self-normalized quantity $\|S_{i}^{(t-1)}\|_{V_{i}^{(t-1),-1}}$ (which can be applied thanks to the conditional $L_{\mu_{i}}$-subgaussianity of $\eta_{i}^{(t)}$ established in Lemma 2.6333In Russac et al. (2019), the subgaussianity of the error term is the sole assumption required to prove their Proposition 5, which in turn implies Proposition 7.), which establishes that
@@
We use the Elliptical Potential Lemma in Carpentier et al. (2020) Proposition 1 (by replacing $p\leftarrow 1$ and $\lambda\leftarrow\frac{\lambda}{c_{\mu_{j}}}$):